# WebDevJudge: Evaluating (M)LLMs as Critiques for Web Development Quality

### v1

@@ 4.2 Different Evaluation Guides and Observation Forms
Model Single Answer Grading Pairwise Comparison Image Only Code Only Both Image Only Code Only Both Claude-4-sonnet 54.43 ($\downarrow$ 1.07) 58.87 ($\uparrow$ 3.37) 55.50 57.34 ($\downarrow$ 8.72) 64.83 ($\downarrow$ 1.23) 66.06 GPT-4.1 53.98 ($\downarrow$ 4.89) 59.94 ($\uparrow$ 1.07) 58.87 57.34 ($\downarrow$ 8.72) 64.98 ($\downarrow$ 1.08) 66.06 GPT-4o 50.31 ($\downarrow$ 5.65) 56.57 ($\uparrow$ 0.61) 55.96 59.48 ($\downarrow$ 4.89) 63.46 ($\downarrow$ 0.91) 64.37 Gemini-2.5-pro 52.29 ($\uparrow$ 3.82) 48.93 ($\uparrow$ 0.46) 48.47 59.17 ($\downarrow$ 3.06) 60.55 ($\downarrow$ 1.68) 62.23 Qwen-2.5-VL-72B 50.00 ($\uparrow$ 1.07) 49.39 ($\uparrow$ 0.46) 48.93 57.19 ($\downarrow$ 6.42) 62.39 ($\downarrow$ 1.22) 63.61
@@ 4.3 Error Analysis
Claude-4-sonnet 89.6 87.9 5.7 3.7 4.7 8.4 0.9 1.38
GPT-4.1 83.3 85.2 0.9 3.0 15.8 11.8 0.8 0.46
DeepSeek-V3-0324 84.1 83.5 11.6 7.3 4.3 9.2 0.5 1.38
@@
Prior studies (Ye et al., 2025) have revealed that LLM-as-a-judge exhibits various forms of bias, such as positional and verbosity biases. We focus specifically on positional bias. Despite explicit instructions to ignore order and remain objective, models still exhibit a systematic preference for responses in a specific position, as shown in Table 5. One might hypothesize that this bias emerges primarily in ambiguous cases where the two options are of comparable quality. However, our analysis of the label distribution for instances with inconsistent predictions reveals that the proportion of ties is not higher than that of wins or losses. This finding suggests that positional bias is not merely an artifact of ambiguity but rather an inherent deficiency in the models, and instruction alone is insufficient to eliminate these deeply embedded inductive biases.
Figure 4: An example illustrating the failure to recognize functional equivalence. The webpage element for ‘Presentation’ serves the same purpose as the required ‘Demonstration’ rating.
@@
We further investigate the evaluators’ ability to accurately verify task fulfillment. Existing benchmarks such as AgentBench (Lù et al., 2025) are too general to assess web-specific judgment. To address this, we construct WebDevJudge-Unit, a targeted dataset of 502 test cases. Each instance consists of web code, a specific verification task, an expected result, and a label indicating whether the task is feasible. We evaluate both LLM-based (code-only) and agent-based (interaction-driven) evaluators on this dataset. From the results in Table 6, we identify complementary weaknesses: (1) For LLM evaluators, these models achieve moderate recall but suffer from low precision. While they can often identify relevant code segments, they are unable to verify actual execution outcomes, leading to false positives when the code appears relevant but does not correctly implement the desired functionality. (2) For agentic evaluators, they exhibit higher precision but lower recall. They are effective when they successfully execute a test plan, but sometimes fail to complete tasks due to limitations in navigation or state interpretation. Consequently, they incorrectly label feasible tasks as infeasible due to their own operational failures rather than actual shortcomings in the web implementation. This divergence highlights a fundamental trade-off: static code analysis lacks execution grounding, while interactive agents are constrained by their own operational reliability. An ideal evaluator would combine the comprehensive coverage of code-aware reasoning with the grounded verification of interactive testing.
Our analysis reveals that the core limitation of LLM-as-a-judge lies in fundamental deficiencies in calibration capability. Lacking this calibration, LLMs struggle to map abstract quality dimensions onto discrete scores and verifiable rubrics. Improving judge performance will require addressing these core competency gaps rather than merely refining evaluation protocols.
5 Conclusion and Future Work
@@
We provide detailed descriptions of our methods, datasets, and evaluation metrics in the main text and appendix to ensure transparency and reproducibility. Our benchmark construction process for WebDevJudge  is detailed in Section 3 and Appendix A. The experimental setup, including the models, evaluation paradigms, and specific protocols, is described in Section 4 and Appendix B. We also provide a complete description of the WebDevJudge-Unit dataset in Appendix C. Code and scripts are available to replicate the empirical results.
Appendix A Details in WebDevJudge Construction
A.1 Data Collection and Filtering Pipeline
Our data collection process commences with the webdev-arena-preference-10k dataset (Vichare et al., 2025; Chiang et al., 2024), which contains 10,501 user queries, each paired with two web implementations and a user-provided preference label. To ensure the quality and suitability of this data for our benchmark, we implemented a rigorous two-stage filtering pipeline.
@@ A.1 Data Collection and Filtering Pipeline
The raw queries were first processed to address quality issues. We began by removing all verbatim duplicate queries to eliminate redundancy. Subsequently, we employed gemini-2.5-pro to screen for and exclude any queries containing harmful, offensive, or nonsensical content, based on a predefined set of safety and clarity instructions. We then exclude queries with minimal interaction requirements (score >> 8, max = 10) and queries with unclear intentions (score << 3, max = 5). The prompts used for the safety filter, interaction filter, and intention filter are presented below.
@@
        "reason": "A clone of WhatsApp Chat App",
        "score": "4"
Environment-based Filtering
This comprehensive filtering process yielded a final set of 1,713 high-quality instances for subsequent annotation. While the original dataset includes user-provided preference labels, we identified several factors that render them unreliable for rigorous evaluation. Primarily, the labels are susceptible to a high degree of subjectivity and variance in evaluation criteria among individual users. The single-pass nature of the crowdsourced annotations also lacks the verification necessary for robust benchmark data. Furthermore, the dataset includes a “tie (bothbad)” category, which introduces ambiguity. The subjective definition of “bad” can lead to inconsistent labeling, potentially masking instances where one implementation, though imperfect, is demonstrably superior to the other. To illustrate these inconsistencies, we provide examples of problematic cases from the original dataset in Table 7.
@@
A.2 Rubric Annotation Pipeline
@@
The prompt used to generate these rubric trees is provided below:
@@
Upon manual review, the LLM-generated rubrics exhibited both strengths and weaknesses. For queries that referenced real-world applications, the rubrics were often of high quality, sometimes surpassing human-authored versions in detail. However, for vague or overly general queries, which are common on crowdsourcing platforms, the generated rubrics sometimes included criteria that were either too specific or too broad, reflecting the ambiguity of the original request. We also provide the statistics of the rubric trees; the details are shown in Table 8. From the statistics, especially the average number of leaf nodes, we can see that the LLM-generated rubrics cover a wide range of criteria, which is a good sign for the evaluation of web development tasks.
Table 8: Statistics of the generated rubric trees.
Metric Intention Static Dynamic Whole
Average height 2.0 3.9 3.3 5.0
Average number of leaf nodes 3.6 15.6 10.6 29.9
To ensure consistent application of these rubrics, we established a set of clear annotation guidelines for the expert annotators, as detailed in the following:
This structured pipeline, combining LLM-generated rubrics with clear human oversight and guidelines, proved highly effective. The final inter-annotator agreement rate reached 89.7%, confirming the consistency and objectivity of our annotation process.
A.3 Data Statistics
@@
Appendix B Experimental Details
@@
B.1 Models And Hyperparameters
Our experiments utilize a combination of open-source and commercial models. Open-source models were deployed using vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024). Commercial models were accessed via their respective APIs, including Azure [^1] https://learn.microsoft.com/en-us/azure/cognitive-services/openai/reference and Vertex AI [^2] https://cloud.google.com/vertex-ai/generative-ai/docs/learn/overview. To ensure reproducibility, we set the temperature to 0.0 for all models, with the exception of claude-3.7-sonnet and claude-4-sonnet, for which the temperature was set to 1.0 to enable their “thinking” mode. For brevity, the models referred to as DeepSeek-V3 and DeepSeek-R1 in the main paper correspond to DeepSeek-V3-0324 and DeepSeek-R1-0528, respectively.
B.2 Evaluation Protocols and Metrics
@@
Motivated by the work of Bian et al. (2025), we designed a multi-level Likert scale for evaluating web development tasks. The scale is based on the international standard for software quality assessment (ISO/IEC/IEEE, 2022) and has been specifically adapted for our dataset. The details of the Likert scale are shown below:
@@ 1. Functional Correctness and Completeness
For single-implementation evaluation, the model assigns a score from 1 (lowest) to 5 (highest) to each sub-dimension. The final score is the sum of all sub-dimension scores. For pairwise comparison, both implementations are presented to the model simultaneously, which allows it to assign scores based on their relative merits. The final preference is determined by the score difference between the two implementations. A preference is declared for the higher-scoring implementation if the score difference exceeds a threshold of 2; otherwise, the outcome is considered a tie.
Rubric
In the single-answer grading setting, the model assigns a binary label (pass or fail) to each leaf node of the rubric. We then compute a pass rate of the leaf node for each of the three primary dimensions: intention, static, and dynamic. The final score is the weighted sum of these pass rates. For pairwise comparison, the model evaluates the two implementations against each leaf node, determining which is superior or if they are tied. This yields a win rate for each implementation of the leaf nodes across the three dimensions. A final score for each is calculated as a weighted sum of these win rates. Preference is awarded to the implementation with the higher final score. In our implementation, all dimension weights are set to 1.
Direct
For direct evaluation, instead of providing any criteria, we instruct the model to directly output its preference (i.e., Wa, Wb, or tie). The prompt we used is shown below.
Prompt for direct comparison You are tasked with comparing two React code snippets (Model A and Model B) based on the user’s query. The input will contain: - User Query: The question or concern the user has. - Answer from Model A: The output of Model A. - Answer from Model B: The output of Model B. Your output should be in the following JSON format: ‘‘‘ { “reason”: “Detailed explanation for why one model is better, based on the user’s query”, “winner”: “‘model_a’, ‘model_b’, or ‘tie”’ } ‘‘‘ - reason: A detailed explanation of why one model’s answer is better than the other, based on the user’s query. Focus on the quality of the responses and how well they address the user’s concerns. - winner: Select the winner from the following options: - “model_a”: If Model A is better. - “model_b”: If Model B is better. - “tie”: If both models are equally good or both provide unsatisfactory answers. Ensure that you thoroughly evaluate the responses before selecting the winner and providing the reason.
B.3 Agentic Workflow Implementation
This section details our agentic workflow implementation. To ensure comparability with our other evaluation methods, the workflow’s planner adopts the LLM-generated rubric tree to guide the executor’s verification actions. The executor is UI-TARS-1.5 (Seed, 2025), one of the state-of-the-art end-to-end GUI agents operating on a ReAct-style paradigm (Yao et al., 2023), where it first generates a thought and then a corresponding action. These actions are converted into executable pyautogui [^3] https://pyautogui.readthedocs.io/en/latest/ code. We utilize the official UI-TARS-1.5 action space and have designed specific prompting strategies for each rubric dimension. For the Static dimension, the rubric tree is converted into a list of elements; the agent navigates the page to find these elements and returns a list of those present. For the Dynamic and Intention dimensions, each rubric leaf node becomes a task for the agent to complete. Upon reaching the maximum number of steps, as detailed in Table 9, the agent provides a conclusion on the task’s outcome. Finally, the summarizer calculates a pass rate for each dimension based on the agent’s findings, and the final score is the weighted sum of these pass rates, mirroring the single-implementation rubric evaluation.

### v2

@@ 4.2 Different Evaluation Guides and Observation Forms
Model Single Answer Grading Pairwise Comparison Image Only Code Only Both Image Only Code Only Both Claude-4-sonnet 55.66 ($\downarrow$ 3.51) 61.77 ($\uparrow$ 2.60) 59.17 59.48 ($\downarrow$ 10.7) 67.58 ($\downarrow$ 2.60) 70.18 GPT-4.1 54.59 ($\downarrow$ 6.27) 62.69 ($\uparrow$ 1.83) 60.86 60.55 ($\downarrow$ 9.79) 69.57 ($\downarrow$ 0.77) 70.34 GPT-4o 53.36 ($\downarrow$ 4.29) 56.27 ($\downarrow$ 1.38) 57.65 59.79 ($\downarrow$ 7.95) 67.43 ($\downarrow$ 0.31) 67.74 Gemini-2.5-pro 53.36 ($\uparrow$ 3.05) 50.15 ($\downarrow$ 0.16) 50.31 62.08 ($\downarrow$ 3.67) 64.53 ($\downarrow$ 1.22) 65.75 Qwen-2.5-VL-72B 51.53 ($\uparrow$ 1.53) 50.92 ($\uparrow$ 0.92) 50.00 59.63 ($\downarrow$ 6.73) 64.68 ($\downarrow$ 1.68) 66.36
@@ 4.3 Error Analysis
Claude-4-sonnet 89.6 87.9 5.7 3.7 4.7 8.4 0.6 1.53
GPT-4.1 83.3 85.2 0.9 3.0 15.8 11.8 0.3 1.07
DeepSeek-V3-0324 84.1 83.5 11.6 7.3 4.3 9.2 0.5 1.23
@@
Prior studies (Ye et al., 2025) have revealed that LLM-as-a-judge exhibits various forms of bias, such as positional and verbosity biases. We focus specifically on positional bias. Despite explicit instructions to ignore order and remain objective, models still exhibit a systematic preference for responses in a specific position, as shown in Table 5. One might hypothesize that this bias emerges primarily in ambiguous cases where the two options are of comparable quality. However, our analysis of the label distribution for instances with inconsistent predictions reveals that the proportion of ties is not higher than that of wins or losses. This finding suggests that positional bias is not merely an artifact of ambiguity but rather an inherent deficiency in the models, and instruction alone is insufficient to eliminate these deeply embedded inductive biases. Rather than employing debiasing techniques such as swapping, we choose the prompting method to reflect the models’ authentic single-pass evaluation capability. Any inherent bias (such as position bias) is considered an intrinsic flaw of the model’s judgment ability. A further comparison with the debiasing technique is provided in Appendix E.1.
Figure 4: An example illustrating the failure to recognize functional equivalence. The webpage element for “Presentation” serves the same purpose as the required “Demonstration” rating. Detailed examples can be found in Appendx F.2.
@@
We further investigate the evaluators’ ability to accurately verify task fulfillment. Existing benchmarks such as AgentRewardBench (Lù et al., 2025) are too general to assess web-specific judgment. To address this, we construct WebDevJudge-Unit, a targeted dataset of 502 test cases. Each instance consists of web code, a specific verification task, an expected result, and a label indicating whether the task is feasible. We evaluate both LLM-based (code-only) and agent-based (interaction-driven) evaluators on this dataset. From the results in Table 6, we identify complementary weaknesses: (1) For LLM evaluators, these models achieve moderate recall but suffer from low precision. While they can often identify relevant code segments, they are unable to verify actual execution outcomes, leading to false positives when the code appears relevant but does not correctly implement the desired functionality. (2) For agentic evaluators, they exhibit higher precision but lower recall. They are effective when they successfully execute a test plan, but sometimes fail to complete tasks due to limitations in navigation or state interpretation. Consequently, they incorrectly label feasible tasks as infeasible due to their own operational failures rather than actual shortcomings in the web implementation. This divergence highlights a fundamental trade-off: static code analysis lacks execution grounding, while interactive agents are constrained by their own operational reliability. An ideal evaluator would combine the comprehensive coverage of code-aware reasoning with the grounded verification of interactive testing. To further investigate this, we design a gated model ensemble strategy to leverage these strengths in the single-comparison for rubric. Specifically, we adopted the results of the LLMs for the intention and static tasks. For the dynamic tasks, we implemented the following logical expression:
$Res_{\text{dynamic}}=Agent\vee LLM$ (2)
This formulation is motivated by the performance trade-off between the constituent components: the Agent provides high-precision filtering for dynamic elements, whereas LLMs ensure higher recall. This hybrid approach optimizes the overall detection capability by balancing exactness with coverage. We selected GPT-4.1, DeepSeek-V3-0324, and UI-TARS 1.5 for this ensemble. The results are shown in Table 7. These results demonstrate that combining static LLM with agent-based dynamic verification can enhance the overall evaluation agreement.
Table 7: Performance comparison between single LLM and the agent-integrated ensemble.
Model Agreement gpt-4.1 65.0     +UI-TARS-1.5 66.2 DeepSeek-V3-0324 66.3     +UI-TARS-1.5 67.6
Our analysis reveals that the limitation of LLM-as-a-judge lies in fundamental deficiencies in calibration capability. Lacking this calibration, LLMs struggle to map abstract quality dimensions onto discrete scores and verifiable rubrics. Improving judge performance will require addressing these competency gaps rather than merely refining evaluation protocols.
5 Conclusion
@@
Acknowledgments
Some authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), and the GRF (16205322) from RGC of Hong Kong, SAR, China. We also thank the support from Tencent AI Lab.
@@ Acknowledgments
We provide detailed descriptions of our methods, datasets, and evaluation metrics in the main text and appendix to ensure transparency and reproducibility. Our benchmark construction process for WebDevJudge  is detailed in Section 3 and Appendix B. The experimental setup, including the models, evaluation paradigms, and specific protocols, is described in Section 4 and Appendix C. We also provide a complete description of the WebDevJudge-Unit dataset in Appendix D. Code and scripts are provided in the supplementary materials to replicate the empirical results.
Appendix A Limitations of WebDevJudge
While WebDevJudge establishes a systematic framework for benchmarking LLM preference prediction, we recognize several constraints that reflect the inherent complexity of modeling human judgment. The core of preference modeling lies in its inherent subjectivity. Although we provided annotators with comprehensive guidelines to standardize evaluations, personal preferences are difficult to fully decouple from objective quality. Because our annotation pool is limited in size, the resulting labels may reflect the collective biases of a specific demographic or professional group rather than a universal human consensus, especially for ambiguous cases. The current scale of WebDevJudge is relatively focused. However, we have retained the broader set of filtered samples, providing a pathway for future data augmentation.
Appendix B Details in WebDevJudge Construction
B.1 Data Collection and Filtering Pipeline
Our data collection process commences with the webdev-arena-preference-10k dataset (Vichare et al., 2025; Chiang et al., 2024), which contains 10,501 user queries, each paired with two web implementations and a user-provided preference label. To ensure the quality and suitability of this data for our benchmark, we implemented a rigorous two-stage filtering pipeline. The overview of the filtering pipeline can be seen in Table 8.
Table 8: Overview of the filtering pipeline, including the number of instances before and after filtering, and the purpose of each filtering stage.
Stage and criterion # before # after Purpose
Query-based: verbatim-identical duplicate 10,501 6,730 Remove redundancy while mataining the original distribution.
Query-based: intention, interaction and safety 6,730 2,460 Remove harmful, offensive, or nonsensical content, and queries with minimal interaction requirements and unclear intentions.
Env-based: deployment failure checking via screenshot 2,460 1,814 Remove instances with deployment failures.
Env-based: deployment failure checking via status code 1,814 1,713 Remove instances with deployment failures.
Sampling 1,713 700 Sample 700 instances regarding the cost-effectiveness.
Manual filtering during annotation 700 654 Filter out instances with harmful content and deployment failures.
@@ B.1 Data Collection and Filtering Pipeline
The raw queries were first processed to address quality issues. We began by removing all verbatim duplicate queries to eliminate redundancy. Subsequently, we employed gemini-2.5-pro to screen for and exclude any queries containing harmful, offensive, or nonsensical content, based on a predefined set of safety and clarity instructions. We then exclude queries with minimal interaction requirements (score << 8, max = 10) and queries with unclear intentions (score << 3, max = 5). The prompts used for filter stages are presented below.
Environment-based Filtering
This comprehensive filtering process yielded a final set of 1,713 high-quality instances for subsequent annotation. We further sampled 700 instances from the final set for annotation. During annotation, we filtered out instances with harmful content and deployment failures manually. While the original dataset includes user-provided preference labels, we identified several factors that render them unreliable for rigorous evaluation. Primarily, the labels are susceptible to a high degree of subjectivity and variance in evaluation criteria among individual users. The single-pass nature of the crowdsourced annotations also lacks the verification necessary for robust benchmark data. Furthermore, the dataset includes a “tie (bothbad)” category, which introduces ambiguity. The subjective definition of “bad” can lead to inconsistent labeling, potentially masking instances where one implementation, though imperfect, is demonstrably superior to the other. To illustrate these inconsistencies, we provide examples of problematic cases from the original dataset in Table 9.
@@
[{"reason": "A clone of WhatsApp Chat App", "score": "4"},]
@@
B.2 Rubric Annotation Pipeline
@@
The prompt used to generate these rubric trees is provided below.
Upon manual review, the LLM-generated rubrics exhibited both strengths and weaknesses. For queries that referenced real-world applications, the rubrics were often of high quality, sometimes surpassing human-authored versions in detail. However, for some general queries, the generated rubrics sometimes included criteria that were either too specific or too broad, reflecting the ambiguity of the original request. We present several examples of the rubric tree structure, encompassing the manually-curated example provided in the one-shot rubric generation, alongside instances of both relatively high and low quality LLM-generated rubrics, as illustrated in Figure 9 in Appendix F.1.
Table 10: Statistics of the generated rubric trees.
Metric Intention Static Dynamic Whole
Average height 2.0 3.9 3.3 5.0
Average number of leaf nodes 3.6 15.6 10.6 29.9
We also provide the statistics of the rubric trees; the details are shown in Table 10. From the statistics, especially the average number of leaf nodes, we can see that the LLM-generated rubrics cover a wide range of criteria, which is a good sign for the evaluation of web development tasks. To ensure consistent application of these rubrics, we established a set of clear annotation guidelines for the expert annotators, as detailed in the following. This structured pipeline, combining LLM-generated rubrics with clear human oversight and guidelines, proved highly effective.
@@
B.3 Data Statistics
@@
Appendix C Experimental Details
@@
C.1 Models And Hyperparameters
Our experiments utilize a combination of open-source and commercial models. Open-source models were deployed using vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024). Commercial models were accessed via their respective APIs, including Azure [^1] https://learn.microsoft.com/en-us/azure/cognitive-services/openai/reference and Vertex AI [^2] https://cloud.google.com/vertex-ai/generative-ai/docs/learn/overview. To ensure reproducibility, we set the temperature to 0.0 for all non-reasoning models. For reasoning models, the temperature was set accordingly. For brevity, the models referred to as DeepSeek-V3 and DeepSeek-R1 in the main paper correspond to DeepSeek-V3-0324 and DeepSeek-R1-0528, respectively.
C.2 Evaluation Protocols and Metrics
@@
Motivated by the work of Bian et al. (2025), we designed a multi-level Likert scale for evaluating web development tasks. The scale is based on the international standard for software quality assessment (ISO/IEC/IEEE, 2022) and has been specifically adapted for our dataset. The details of the Likert scale are shown in “Dimensions of the Likert Scale”.
For single-implementation evaluation, the model assigns a score from 1 (lowest) to 5 (highest) to each sub-dimension. The final score is the sum of all sub-dimension scores. For pairwise comparison, both implementations are presented to the model simultaneously, which allows it to assign scores based on their relative merits. The final preference is determined by the score difference between the two implementations. A preference is declared for the higher-scoring implementation if the score difference exceeds a threshold of 1; otherwise, the outcome is considered a tie.
Rubric
In the single-answer grading setting, the model assigns a binary label (pass or fail) to each leaf node of the rubric. We then compute a pass rate of the leaf node for each of the three primary dimensions: intention, static, and dynamic. The final score is the weighted sum of these pass rates. For pairwise comparison, the model evaluates the two implementations against each leaf node, determining which is superior or if they are tied. This yields a win rate for each implementation of the leaf nodes across the three dimensions. A final score for each is calculated as a weighted sum of these win rates. Preference is awarded to the implementation with the higher final score. In our implementation, dimension weights for intention, static, dynamic are set to 1, 3, 2, respectively.
Direct
For direct evaluation, instead of providing any criteria, we instruct the model to directly output its preference (i.e., Wa, Wb, or tie).
@@ 1. Functional Correctness and Completeness
C.3 Agentic Workflow Implementation
This section details our agentic workflow implementation, as shown in Table 11. To ensure comparability with our other evaluation methods, the workflow’s planner adopts the LLM-generated rubric tree to guide the executor’s verification actions. The executor is UI-TARS-1.5 (Seed, 2025), one of the state-of-the-art end-to-end GUI agents operating on a ReAct-style paradigm (Yao et al., 2023), where it first generates a thought and then a corresponding action. These actions are converted into executable pyautogui [^3] https://pyautogui.readthedocs.io/en/latest/ code. We utilize the official UI-TARS-1.5 action space and have designed specific prompting strategies for each rubric dimension. For the Static dimension, the rubric tree is converted into a list of elements; the agent navigates the page to find these elements and returns a list of those present. For the Dynamic and Intention dimensions, each rubric leaf node becomes a task for the agent to complete. Upon reaching the maximum number of steps, the agent provides a conclusion on the task’s outcome. Finally, the summarizer calculates a pass rate for each dimension based on the agent’s findings, and the final score is the weighted sum of these pass rates, mirroring the single-implementation rubric evaluation.